# A Bayes Factor Framework for Unified Parameter Estimation and Hypothesis Testing

### v3

@@ 2.4 Choice of the prior
As the previous example showed, the prior assigned to the parameters under the alternative has a substantial impact on SC inference. This ‘sensitivity’ of Bayes factors to prior distributions enables data analysts to accurately quantify the support of parameter values over informative alternative hypotheses when they are available, but poses a challenge in their absence (Kass and Raftery, , 1995). Various approaches have been proposed to deal with this issue, for example, ‘default’ or ‘objective’ prior distributions (Bayarri et al., , 2012; Consonni et al., , 2018), reverse-Bayes analysis (Held et al., , 2022), prior elicitation (O’Hagan, , 2019), or sensitivity analysis (Franck and Gramacy, , 2019), all with advantages and disadvantages. Here we will not reiterate general considerations on prior specification for Bayes factors (see e.g., Section 5 in Kass and Raftery, , 1995) but focus on specific considerations related to SCs.
@@
As in other Bayes factor applications, SCs are only unambiguously defined if priors for focus parameters are proper under the alternative $H_{1}$ (i.e., integrate to one), whereas priors for nuisance parameters may be improper as long as the same prior is assigned under both the null $H_{0}$ and the alternative $H_{1}$ so that arbitrary constants cancel out. A general distinction can be made between global priors, which do not depend on the value of $\theta_{0}$ under the null hypothesis and local priors, which do. In the latter case, the interpretation of the SC is more intricate, since for each parameter value the SC quantifies the support over a different alternative. For a more natural interpretation, global priors may hence be preferred over local priors. At the same time, local priors correspond to the typical use of ‘default’ Bayes factors, which is to center the prior around $\theta_{0}$, and as such may be preferred in the same situations where default Bayes factors would be used.
@@
Finally, it is usually advisable to report sensitivity analyses for plausible ranges of priors, to assess the robustness of the conclusions. A convenient visual sensitivity analysis is, for example, to plot different SCs resulting from different prior specifications, as shown in Figure 2. One can go a step further and use a ‘reverse-Bayes‘ approach (Good, , 1950; Held et al., , 2022), which involves systematically determining the prior that represents the tipping point and changes the conclusions of the analysis. Data analysts can then reason about whether or not such a prior is plausible in the light of external knowledge and data.
@@ 2.6 Asymptotic behaviour of the support curve
As a concrete example where the distribution of the SC can be derived in closed-form, consider again inference about a normal mean based on data $Y\mid\theta\sim\operatorname{N}(\theta,\kappa^{2}/n)$, where $\kappa^{2}$ denotes a unit-variance and $n$ the sample size. The logarithm of the SC based on a normal prior $\theta\mid H_{1}\sim\operatorname{N}(m,v)$ can then be written as
$\log\operatorname{BF}_{01}(Y;\theta_{0})$ $=\frac{1}{2}\left[\log\left(1+\frac{n\,v}{\kappa^{2}}\right)+\frac{(\theta_{0}-m)^{2}}{v}-\left\{Y-\frac{(\theta_{0}-m)\kappa^{2}}{n\,v}-\theta_{0}\right\}^{2}\frac{v\,n}{\kappa^{2}(v+\kappa^{2}/n)}\right].$ (3)
Hence, when the data are generated from $Y\mid\theta_{*}\sim\operatorname{N}(\theta_{*},\kappa^{2}/n)$ with true mean $\theta_{*}$, we have that
$\left\{Y-\frac{(\theta_{0}-m)\kappa^{2}}{n\,v}-\theta_{0}\right\}^{2}\frac{n}{\kappa^{2}}\sim\chi^{2}_{1,\lambda}$
@@
$\Pr\{\operatorname{BF}_{01}(Y;\theta_{0})\leq\gamma\mid\theta_{*}\}=1-\Pr(\chi_{1,\lambda}^{2}\leq X)$
@@
Figure 3: Distribution of the SC for different sample sizes. A data model $Y\mid\theta\sim\operatorname{N}(\theta,\kappa^{2}/n)$ is assumed and data are generated from a true mean $\theta_{*}=0$ and unit-variance $\kappa^{2}=4$. The SC is based on a local normal prior $\theta\mid H_{1}\sim\operatorname{N}(\theta_{0},v=4)$ assigned to $\theta$ under the alternative.
@@ 3.1 Maximum integrated likelihood
The same MEE is obtained when the prior $p(\mu,\sigma^{2}\mid H_{1})$ does not depend on the value of the variance under $H_{0}$ as the denominator of the Bayes factor is simply a multiplicative factor that does not change its maximum. This shows that REML and MIL estimates can also be motivated from a Bayesian evidence perspective which complements the well-established connections between REML estimation and marginal posterior estimation based on flat priors for the nuisance parameters (Harville, , 1974; Laird and Ware, , 1982). It is reassuring that different methods produce the same estimate in these situations. However, the important difference between these methods is the motivation and interpretation of the resulting estimate – the MEE represents a natural estimate for $\theta$ because it is the parameter value for which the data provide the most evidence over an alternative hypothesis, while the (integrated) maximum likelihood estimate is defined without reference to alternatives.
@@ 3.3 Frequentist inference
Another connection between frequentist and SC inference is given by the ‘universal bound’ (Kerridge, , 1963; Robbins, , 1970; Royall, , 1997; Sanborn and Hills, , 2013), which bounds the frequentist probability of obtaining misleading Bayesian evidence. That is, when data are generated under $H_{0}:\theta=\theta_{0}$, the probability of obtaining a Bayes factor $\operatorname{BF}_{01}$ less than $k<1$ is at most $k$ for any prior under the alternative
@@
If there are nuisance parameters, the bound holds only marginalized over the prior of the nuisance parameters. For the bound to hold in a strict sense (i.e., for every possible value of the nuisance parameter), special priors must be assigned to them (Hendriksen et al., , 2021; Grünwald et al., , 2024).
The universal bound can thus be used to transform SCs into conservative P-values and confidence sets, e.g., a $k=1/20$ support set obtained from a SC corresponds to a 95% conservative confidence set and $p=\min\{\operatorname{BF}_{01},1\}$ corresponds to a conservative $P$-value. Remarkably, the bound holds without adjustment even when the data collection is continuously monitored and stopped as soon as evidence against $H_{0}$ is found (Robbins, , 1970). However, it is important to note that P-values and confidence sets obtained in this way are usually much more conservative than ordinary ones which are calibrated to have exact type I error rate and coverage, respectively. Finally, if the data model is misspecified, the bound is obviously invalid.
@@ 3.4 Bayesian inference
The SC can, under certain conditions, be transformed into a Bayesian posterior distribution. Specifically, assuming that the priors for the nuisance parameters satisfy $p(\psi\mid H_{0})=p(\psi\mid\theta=\theta_{0},H_{1})$, the Bayes factor can be represented as the ratio of marginal posterior to prior density evaluated at the tested parameter value (known as Savage-Dickey density ratio, see e.g., Dickey, , 1971; Verdinelli and Wasserman, , 1995; Wagenmakers et al., , 2010). Hence, the posterior can be obtained by multiplying the SC with the prior
@@
The Savage-Dickey density ratio (5) also provides a convenient way to compute SCs: One of the many programs for computing Bayesian posterior distributions, such as Stan (Carpenter et al., , 2017) or INLA (Rue et al., , 2009), can be used to compute a posterior density, which can then be divided by the prior density to obtain a SC. The caveat is again that this only works for global priors under the alternative and with the same prior assigned to the nuisance parameters under the null and alternative.
@@ 4.2 Meta-analysis
$\theta_{i}\mid\theta,\tau^{2}$ $\sim\operatorname{N}(\theta,\tau^{2}).$
Marginalized over the participant specific probabilities, the distribution of an estimate is then
@@
There are two unknown parameters, $\theta$ and $\tau$. The mean $\theta$ quantifies the average probability across participants, while the heterogeneity standard deviation $\tau$ quantifies the heterogeneity of these probabilities. The Bayes factor for testing $H_{0}\colon\theta=\theta_{0},\tau=\tau_{0}$ against $H_{1}\colon\theta\neq\theta_{0},\tau\neq\tau_{0}$ is then given by
$\operatorname{BF}_{01}(y_{1},\dots,y_{n};\theta_{0},\tau_{0})=\frac{\prod_{i}^{n}\operatorname{N}(y_{i}\mid\theta_{0},\tau_{0}^{2})}{\int_{0}^{\infty}\int_{-\infty}^{+\infty}\prod_{i}^{n}\operatorname{N}(y_{i}\mid\theta,\tau^{2})\,p(\theta,\tau\mid H_{1})\,\mathrm{d}\theta\,\mathrm{d}\tau}$
@@ 4.3 Replication studies
$y_{r}\pm\sigma_{r}\sqrt{\log\left(1+\frac{\sigma^{2}_{o}}{\sigma^{2}_{r}}\right)+\frac{(y_{r}-y_{0})^{2}}{\sigma^{2}_{r}+\sigma^{2}_{o}}-\log k^{2}}.$
@@
We will now demonstrate application of the replication SC by reanalyzing data from the replication project by Wagenmakers et al., (2016). This project attempted to replicate the original study from Strack et al., (1988) across 17 different study sites. The original study tested the so-called ‘facial feedback hypothesis’ and found that participants gave higher funniness ratings to cartoons if they were smiling as opposed to showing discontent (estimated mean difference of $0.82$ units on a 10-point Likert scale, with 95% confidence interval from $-0.05$ to $1.69$). The replications were conducted across 17 different study sites, each producing a mean difference effect estimate and confidence interval. In contrast to the original study, the pooled replication mean difference was very close to zero (estimated mean difference of $0.03$ with 95% confidence interval from $-0.11$ to $0.16$). We will now assess the replicability of the original finding using replication Bayes factors and the corresponding support curves.
@@
Figure 6: Support curves with maximum evidence estimate and $k=1$ support interval (top) and posterior distribution with posterior mode and 95% highest posterior density credible interval (bottom) for the pooled replication studies from the ‘facial feedback hypothesis’ (Strack et al., , 1988; Wagenmakers et al., , 2016). The original study found an estimated standardized mean difference of $y_{o}=0.82$ with standard error $\sigma_{o}=0.444$ which is used to formulate the prior distribution under the alternative $\theta\mid H_{1}\sim\operatorname{N}(y_{o},\sigma^{2}_{o})$. The replication effect estimates are pooled via inverse variance weighting. The posterior is obtained by multiplying the BFF with the prior density.
@@
The top plot in Figure 6 shows the associated replication SCs, MEE, and $k=1$ support interval. We see from the support curve based on the pooled replications that mean difference values close to zero receive more evidence compared to the prior based on the original study ($\hat{\theta}_{\text{ME}}=0.03$ with $k=1$ support interval from $-0.15$ to $0.21$), all of them being much smaller than the estimated mean difference from the original study 0.82. Furthermore, the support curve at the mean difference of zero indicates strong evidence in favor of no difference over the prior based on the original study ($\mathrm{BF}_{01}=29$). Thus, the replication SC analysis suggests hardly any replicability of the original facial feedback effect.
@@
The bottom plot in Figure 6 illustrates the posterior distributions, conveniently obtained by multiplying the SC by the prior distribution based on the original data since this SC has a Savage-Dickey density ratio representation. We can see that the $k=1$ support intervals from the top plot are given by the set of effect sizes with posterior density larger than the prior density. We can also see that for the pooled replications the support curve inferences mostly align with the posterior inferences, that is, the maximum a posteriori estimate and 95% credible interval are very close to the MEE and $k=1$ support interval. This is because when pooling all the replication data, the data are much more informative than the prior based on the original study and thus overturn it. However, when comparing the individual replications’ support curves and posterior distributions we can see much larger differences. The SCs peak at the corresponding replication effect estimates while the posteriors peak at weighted averages of original and replication estimates. For example, the replication with the smallest effect has a SC with $\hat{\theta}_{\text{ME}}=-0.58$ and with $k=1$ support interval from $-1.62$ to $0.46$, whereas the corresponding posterior has its mode at $\hat{\theta}_{\text{MAP}}=0.09$ with $95\%$ credible interval from $-0.51$ to $0.69$. This apparent discrepancy reflects the fact that SCs do not synthesize the data with the prior, as opposed to the posterior analysis. Since in a replication study the interest is often to challenge the original study, synthesis with the original data is arguably undesirable in the analysis of replication studies. As such, a replication SC analysis, as presented here, can be useful because it can clearly show whether or not there are parameter values that are better supported than the prior based on the original study.

### v4

@@ 2.4 Choice of the prior
While Bayes factors do not depend on the prior probabilities of their contrasted hypothesis, they can be sensitive to the prior distributions specified for model parameters under each hypothesis. Consequently, the parameter priors also have a substantial impact on SC inferences, as the previous example showed. Moreover, unlike posterior distributions, where the influence of the prior diminishes as data accumulate, the sensitivity of Bayes factors to prior distributions persists even with large samples (Kass and Raftery, , 1995).
This sensitivity is not inherently a weakness. It enables data analysts to accurately quantify the support of parameter values over informative alternative hypotheses when they are available (Vanpaemel, , 2010), but poses a challenge in their absence (Liu and Aitkin, , 2008; Williams et al., , 2023). Even when external information exists, translating it into a prior distribution is often nontrivial. For instance, knowing that a parameter lies within a certain range still leaves considerable flexibility in the actual specification of the prior.
Several strategies have been proposed to address these challenges, each with advantages and limitations. First, so-called ‘default’ or ‘objective’ priors aim to yield Bayes factors with desirable properties, such as model selection consistency, without incorporating prior knowledge (Bayarri et al., , 2012; Consonni et al., , 2018). These are often provided as defaults in software for computing Bayes factors, such as the BayesFactor R package (Morey and Rouder, , 2024) or JASP (Love et al., , 2019). Second, prior elicitation techniques can be used to formally translate expert judgments into quantitative priors (O’Hagan, , 2019; Stefan et al., , 2022; Mikkola et al., , 2024). Third, Bayes factor bounds (Berger and Sellke, , 1987; Sellke et al., , 2001; Held and Ott, , 2018) and sensitivity analysis (Kass and Raftery, , 1995; Sinharay and Stern, , 2002; Franck and Gramacy, , 2019) explore how Bayes factors change across a class of priors to assess robustness of conclusions. Finally, reverse-Bayes analyses identify the priors that lead to the Bayes factor crossing a tipping point, allowing analysts to assess the plausibility of such a prior in light of external knowledge (Good, , 1950; Held et al., , 2022). As all of these techniques are applicable to Bayes factors, they are also applicable to SCs.
As in other Bayes factor applications, SCs are only unambiguously defined if priors for focus parameters are proper (i.e., integrate to one) under the alternative $H_{1}$. Assigning improper or very vague priors can lead SCs to diverge to infinity thereby proving overwhelming support for every tested parameter value $\theta_{0}$, which is known as the ‘Jeffreys-Lindley paradox’ (Robert, , 2014; Wagenmakers and Ly, , 2023). Priors for nuisance parameters may be improper as long as the same prior is assigned under both the null $H_{0}$ and the alternative $H_{1}$ so that arbitrary constants cancel out.
@@
A general distinction can be made between global and local priors. Global priors do not depend on the parameter value $\theta_{0}$ of the null hypothesis. In contrast, local priors are defined relative to $\theta_{0}$. For example, a local prior may be centered around $\theta_{0}$ or shifted away from it, as the priors underlying the green and yellow SCs in Figure 2. This results in a different alternative hypothesis for each tested parameter value, making the interpretation of the SC more intricate. For a more natural interpretation, global priors may hence be preferred over local priors. At the same time, local priors correspond to the typical use of default Bayes factors, which is to center the prior around $\theta_{0}$, and as such may be used in the same situations where default Bayes factors would be used.
@@
Given these considerations, SCs, just as Bayes factors, are especially useful if informative priors based on prior data or external knowledege are available. If such priors are lacking and default priors are used, it is advisable to report sensitivity analyses for plausible ranges of priors or reverse-Bayes analysis, to assess the robustness of the conclusions. A convenient visual sensitivity analysis is, for example, to plot different SCs resulting from different prior specifications, as shown in Figure 2.
@@ 2.6 Asymptotic behaviour of the support curve
As a concrete example where the distribution of the SC can be derived in closed-form, consider again inference about a normal mean based on the data mean $\bar{y}$ with distribution $\bar{Y}\mid\theta\sim\operatorname{N}(\theta,\kappa^{2}/n)$, where $\kappa^{2}$ denotes a unit-variance and $n$ the sample size. The logarithm of the SC based on a normal prior $\theta\mid H_{1}\sim\operatorname{N}(m,v)$ can then be written as
$\log\operatorname{BF}_{01}(\bar{Y};\theta_{0})$ $=\frac{1}{2}\left[\log\left(1+\frac{n\,v}{\kappa^{2}}\right)+\frac{(\theta_{0}-m)^{2}}{v}-\left\{\bar{Y}-\frac{(\theta_{0}-m)\kappa^{2}}{n\,v}-\theta_{0}\right\}^{2}\frac{v\,n}{\kappa^{2}(v+\kappa^{2}/n)}\right].$ (3)
Hence, when data means are generated from $\bar{Y}\mid\theta_{*}\sim\operatorname{N}(\theta_{*},\kappa^{2}/n)$ with true mean $\theta_{*}$, we have that
$\left\{\bar{Y}-\frac{(\theta_{0}-m)\kappa^{2}}{n\,v}-\theta_{0}\right\}^{2}\frac{n}{\kappa^{2}}\sim\chi^{2}_{1,\lambda}$
@@
$\Pr\{\operatorname{BF}_{01}(\bar{Y};\theta_{0})\leq\gamma\mid\theta_{*}\}=1-\Pr(\chi_{1,\lambda}^{2}\leq X)$ (4)
@@
Figure 3: The distribution of the SC, summarized by 5% quantile (lower dashed line), 50% quantile (solid line), 95% quantile (upper dashed line), for different sample sizes $n$ (colors). A data model $\bar{Y}\mid\theta\sim\operatorname{N}(\theta,\kappa^{2}/n)$ is assumed with true mean $\theta_{*}=0$ and unit-variance $\kappa^{2}=4$. The SC is based on a local normal prior $\theta\mid H_{1}\sim\operatorname{N}(\theta_{0},v=4)$ assigned to $\theta$ under the alternative. Quantiles were computed by numerically inverting the cumulative distribution function of the Bayes factor (4).
@@ 3.1 Maximum integrated likelihood
The same MEE is obtained when the prior $p(\mu,\sigma^{2}\mid H_{1})$ does not depend on the value of the variance under $H_{0}$ as the denominator of the Bayes factor is simply a multiplicative factor that does not change its maximum. This shows that REML and MIL estimates can also be motivated from a Bayes factor perspective which complements the well-established connections between REML estimation and marginal posterior estimation based on flat priors for the nuisance parameters (Harville, , 1974; Laird and Ware, , 1982). It is reassuring that different methods produce the same estimate in these situations. However, the important difference between these methods is the motivation and interpretation of the resulting estimate – the MEE represents a natural estimate for $\theta$ because it is the parameter value for which the data provide the most evidence over an alternative hypothesis, while the (integrated) maximum likelihood estimate is defined without reference to alternatives.
@@ 3.3 Frequentist inference
Another connection between frequentist and SC inference is given by the ‘universal bound’ (Kerridge, , 1963; Robbins, , 1970; Royall, , 1997; Sanborn and Hills, , 2013), which bounds the frequentist probability of obtaining misleading Bayesian evidence. That is, when data are generated under $H_{0}:\theta=\theta_{0}$, the probability of obtaining misleading evidence against the null at level $k$ (with a $\operatorname{BF}_{01}\leq k<1$) is at most $k$ for any prior under the alternative
@@
This can be shown by noting that the expectation of $\operatorname{BF}_{10}$ under $H_{0}$ is 1 and applying Markov’s inequality (see e.g., Grünwald et al., , 2024). If there are nuisance parameters, the bound holds only marginalized over the prior of the nuisance parameters. For the bound to hold in a strict sense (i.e., for every possible value of the nuisance parameter), special priors must be assigned to them (Hendriksen et al., , 2021; Grünwald et al., , 2024).
The universal bound can thus be used to transform SCs into conservative P-values and confidence sets, e.g., a $k=1/20$ support set obtained from a SC corresponds to a 95% conservative confidence set and $p=\min\{\operatorname{BF}_{01},1\}$ corresponds to a conservative $P$-value. Remarkably, the bound holds without adjustment even in ‘optional stopping’ settings where data collection is continuously monitored and stopped as soon as evidence against $H_{0}$ is found (Robbins, , 1970). However, it is important to note that P-values and confidence sets obtained in this way are usually much more conservative than ordinary ones which are calibrated to have exact type I error rate and coverage, respectively. Finally, if the data model is misspecified, the bound is invalid.
@@ 3.4 Bayesian inference
The SC can, under certain conditions, be transformed into a Bayesian posterior distribution. Specifically, assuming that the priors for the nuisance parameters satisfy
$p(\psi\mid H_{0})=p(\psi\mid\theta=\theta_{0},H_{1}),$ (6)
the Bayes factor can be represented as the ratio of marginal posterior to prior density under $H_{1}$ evaluated at the tested parameter value (known as Savage-Dickey density ratio, see e.g., Dickey, , 1971; Verdinelli and Wasserman, , 1995; Wagenmakers et al., , 2010). Hence, the posterior under $H_{1}$ can be obtained by multiplying the SC with the prior
@@
The Savage-Dickey density ratio (7) also provides a convenient way to compute SCs: One of the many programs for computing Bayesian posterior distributions, such as Stan (Carpenter et al., , 2017) or INLA (Rue et al., , 2009), can be used to compute a posterior density, which can then be divided by the prior density to obtain a SC. The caveat is again that this only works for global priors under the alternative and provided that condition (6) holds for the priors assigned to the nuisance parameters.
@@ 4.2 Meta-analysis
$\theta_{i}\mid\theta,\tau$ $\sim\operatorname{N}(\theta,\tau^{2}).$
While in some applications, inferences about the participant specific probabilities $\theta_{i}$ would also be of interest, these are consider nuisance parameters in this analysis. Consequently, the distribution of an estimate is marginalized over the participant probabilities distribution, leading to the model
@@
There are two unknown parameters left, $\theta$ and $\tau$. The mean $\theta$ quantifies the average probability across participants, while the heterogeneity standard deviation $\tau$ quantifies the heterogeneity of these probabilities. The Bayes factor for testing $H_{0}\colon\theta=\theta_{0},\tau=\tau_{0}$ against $H_{1}\colon\theta\neq\theta_{0},\tau\neq\tau_{0}$ is then given by
$\operatorname{BF}_{01}(y_{1},\dots,y_{n};\theta_{0},\tau_{0})=\frac{\prod_{i}^{n}\operatorname{N}(y_{i}\mid\theta_{0},\sigma^{2}_{i}+\tau_{0}^{2})}{\int_{0}^{\infty}\int_{0}^{1}\prod_{i}^{n}\operatorname{N}(y_{i}\mid\theta,\sigma^{2}_{i}+\tau^{2})\,p(\theta,\tau\mid H_{1})\,\mathrm{d}\theta\,\mathrm{d}\tau}$
@@ 4.3 Replication studies
$y_{r}\pm\sigma_{r}\sqrt{\log\left(1+\frac{\sigma^{2}_{o}}{\sigma^{2}_{r}}\right)+\frac{(y_{r}-y_{o})^{2}}{\sigma^{2}_{r}+\sigma^{2}_{o}}-\log k^{2}}.$
@@
We will now demonstrate application of the replication SC by reanalyzing data from the replication project by Wagenmakers et al., (2016). This project attempted to replicate the original study from Strack et al., (1988) across 17 different study sites. The original study tested the so-called ‘facial feedback hypothesis’ and found that participants gave higher funniness ratings to cartoons if they were smiling as opposed to showing discontent (estimated mean difference of $0.82$ units on a 10-point Likert scale, with 95% confidence interval from $-0.05$ to $1.69$). The replications were conducted across 17 different study sites, each producing a mean difference effect estimate and confidence interval. In contrast to the original study, the pooled replication mean difference was very close to zero (estimated mean difference of $0.03$ with 95% confidence interval from $-0.11$ to $0.16$). We will now assess the replicability of the original finding using replication Bayes factors and the corresponding SCs.
@@
Figure 6: Support curves with maximum evidence estimate and $k=1$ support interval (top) and posterior distribution with posterior mode and 95% highest posterior density credible interval (bottom) for the pooled replication studies from the ‘facial feedback hypothesis’ (Strack et al., , 1988; Wagenmakers et al., , 2016). The original study found an estimated standardized mean difference of $y_{o}=0.82$ with standard error $\sigma_{o}=0.44$ which is used to formulate the prior distribution under the alternative $\theta\mid H_{1}\sim\operatorname{N}(y_{o},\sigma^{2}_{o})$. The replication effect estimates are pooled via inverse variance weighting. The posterior is obtained by multiplying the SC with the prior density.
@@
The top plot in Figure 6 shows the associated replication SCs, MEE, and $k=1$ support interval. We see from the SC based on the pooled replications that mean difference values close to zero receive more evidence compared to the prior based on the original study ($\hat{\theta}_{\text{ME}}=0.03$ with $k=1$ support interval from $-0.15$ to $0.21$), all of them being much smaller than the estimated mean difference from the original study 0.82. Furthermore, the SC at the mean difference of zero indicates strong evidence in favour of no difference over the prior based on the original study ($\mathrm{BF}_{01}=29$). Thus, the replication SC analysis suggests hardly any replicability of the original facial feedback effect.
@@
The bottom plot in Figure 6 illustrates the posterior distributions, conveniently obtained by multiplying the SC by the prior distribution based on the original data since this SC has a Savage-Dickey density ratio representation. We can see that the $k=1$ support intervals from the top plot are given by the set of effect sizes with posterior density larger than the prior density. We can also see that for the pooled replications the SC inferences mostly align with the posterior inferences, that is, the maximum a posteriori estimate and 95% credible interval are very close to the MEE and $k=1$ support interval. This is because when pooling all the replication data, the data are much more informative than the prior based on the original study and thus overturn it. However, when comparing the individual replications’ SCs and posterior distributions we can see much larger differences. The SCs peak at the corresponding replication effect estimates while the posteriors peak at weighted averages of original and replication estimates. For example, the replication with the smallest effect has a SC with $\hat{\theta}_{\text{ME}}=-0.58$ and with $k=1$ support interval from $-1.62$ to $0.46$, whereas the corresponding posterior has its mode at $\hat{\theta}_{\text{MAP}}=0.09$ with $95\%$ credible interval from $-0.51$ to $0.69$. This apparent discrepancy reflects the fact that SCs do not synthesize the data with the prior, as opposed to the posterior analysis. Since in a replication study the interest is often to challenge the original study, synthesis with the original data is arguably undesirable in the analysis of replication studies. As such, a replication SC analysis, as presented here, can be useful because it can clearly show whether or not there are parameter values that are better supported than the prior based on the original study.